# Connectedness of the Gromov boundary of fine curve graphs

### v2

@@ 4 Connectivity properties
In order to demonstrate the connectedness of the Gromov boundary of the fine curve graph, we need only to prove that the spheres in the subgraph $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ enjoys connectivity properties as in the assumption of Proposition 2.11.
For notations, for any positive integer $r>0$, let $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ be the sphere in $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ centred at $o$ of radius $r$, i.e. $\{\alpha\in\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)\colon d^{\dagger}(\alpha,o)=r\}$.
@@
Let $r>0$ be any positive integer and let $\alpha,\beta\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$. If $\alpha$ and $\beta$ are not transverse, then there exists a $\beta^{\prime}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ isotopic to $\beta$ but disjoint from it, which is also transverse to $\alpha$.
@@
In light of Proposition 3.5, we can find a finite subset $P\subset S$ so that $(o,\alpha)$ and $(o,\beta)$ are in minimal positions in $S-P$. Up to a small perturbation as in Lemma 3.7, the curve $\beta$ is isotopic to but disjoint from $\beta^{\prime}$ while $\beta^{\prime}$ is transverse to $\alpha$. Since the perturbation can be taken so that $\beta^{\prime}$ remains in a minimal position with $o$ in $S-P$. By Proposition 3.5, $\beta^{\prime}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$. ∎
@@
For any positive integer $r>0$, the sphere $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ is connected.
@@
Let $\alpha,\beta\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ be distinct non-separating curves. By Lemma 4.1, up to perturbation, we can choose $\beta^{\prime}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ isotopic to but disjoint from $\beta$. Moreover, there exists a finite set $P\subset S$ such that $\beta^{\prime}$, $\alpha$ and $o$ are in minimal positions in $S-P$. By taking a hyperbolic metric on $S-P$ as in Proposition 3.4, we can even assume that $\beta^{\prime}$ (up to perturbation), $\alpha$ and $o$ are geodesics. Moreover, in view of Proposition 3.5, there is an isometric embedding via geodesic representations $\iota\colon\operatorname{\mathcal{NC}}(S-P)\hookrightarrow\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$. Note that $S-P$ is of at least genus $2$ and $1$ puncture, i.e. of high complexity as called in [Wri23]. Then applying [Wri23, Proposition 5.4] and [Wri23, Lemma 2.1] to $\operatorname{\mathcal{NC}}(S-P)$, there is a path $[\alpha]_{S-P}=[\nu_{0}]_{S-P},\ [\nu_{1}]_{S-P},\ \dots,\ [\nu_{n}]_{S-P}=[\beta^{\prime}]_{S-P}$ lying inside of the $r$-sphere centred at $[o]_{S-P}$ in $\operatorname{\mathcal{NC}}(S-P)$. Hence the image under the isometric embedding $\iota$ via geodesic representations, $\alpha=\nu_{0},\ \nu_{1},\ \dots,\ \nu_{n}=\beta^{\prime}$, yields a path connecting $\alpha$ to $\beta^{\prime}$ in $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$. Since $\beta$ and $\beta^{\prime}$ is disjoint, they are adjacent in $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$. Hence the concatenation $\alpha=\nu_{0},\ \nu_{1},\ \dots,\ \nu_{n}=\beta^{\prime},\ \beta$ is a path in $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ connecting the two vertices. ∎
Similar arguments will soon yield Theorem 1.2:
@@
For any two distinct $\beta,\gamma\in\{\alpha\in\operatorname{\mathcal{NC}}^{\dagger}(S)\colon d^{\dagger}(\alpha,o)=r\}$, as in Lemma 4.1, we can similarly perturb them to $\beta^{\prime}$ and $\gamma^{\prime}$ that are isotopic to but disjoint from $\beta$ and $\gamma$ respectively, so that $\{o,\beta^{\prime},\gamma^{\prime}\}$ are transverse and $d^{\dagger}(o,\beta^{\prime})=d^{\dagger}(o,\gamma^{\prime})=r$. Now by Lemma 4.2, we can connect them by path contained in the sphere. ∎
The following lemma shows that there is no “dead ends” in the sense of [BM15] inside of the subgraph $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$, i.e. this subgraph satisfies the assumption (G1) in Proposition 2.11.
Lemma 4.3.
For every adjacent pair $\alpha,\beta\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ there exists a path
$\alpha=\nu_{0},\nu_{1},\nu_{2}=\beta$
in $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ with $\nu_{1}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)$.
Proof.
Since $\alpha$ and $\beta$ are disjoint, take a non-empty finite subset $P\subset S$ so that $P$ is disjoint from $\alpha\cup\beta\cup o$ and that $o,\alpha,\beta$ are in minimal positions. Take a hyperbolic metric on $S-P$ as in Proposition 3.4 so that $\alpha$ and $\beta$ are geodesics. By [Wri23, Proposition 5.4], we can find a geodesic curve $\nu$ with respect to the above hyperbolic metric such that it is disjoint from $\alpha$ and $\beta$ with $d_{\operatorname{\mathcal{NC}}(S-P)}([\nu]_{S-P},[o]_{S-P})=r+1$. Hence it lies in a connected component $\Sigma$ of $S-\alpha\cup\beta$. Now perturb $\nu$ isotopically into $\nu_{1}$ so that it has no bigon with $o\cap\Sigma$ in $\Sigma-P$. This will make sure that $\nu_{1}$ is still disjoint from $\alpha$ and $\beta$, but is in minimal position with $o$ in $S-P$. Note that $[\nu]_{S-P}=[\nu_{1}]_{S-P}$. Hence $d^{\dagger}(o,\nu_{1})=d_{\operatorname{\mathcal{NC}}(S-P)}([\nu_{1}]_{S-P},[o]_{S-P})=r+1$ by Proposition 3.5. ∎
Lemma 4.4.
Let $r>0$ and $\alpha,\alpha^{\prime}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ be isotopic. Then there exists a path on $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ connecting $\alpha$ to $\alpha^{\prime}$, in which every vertices is a curve isotopic to $\alpha$. If $\beta\in\operatorname{\mathcal{NC}}^{\dagger}(S)$ is in addition disjoint from both $\alpha$ and $\alpha^{\prime}$, then so will be the vertices in this path.
Proof.
Let $P\subset S$ be a finite subset so that $(o,\alpha)$ and $(o,\alpha^{\prime})$ are in minimal positions in $S-P$. Up to a small perturbation as in Lemma 3.7, we can take $\alpha^{\prime\prime}$ that is isotopic to but disjoint from $\alpha$ and that is transverse to $\alpha^{\prime}$. Since the perturbation is small, the curve $\alpha^{\prime\prime}$ will remain in minimal position with $o$ on $S-P$ and is disjoint from $\beta$. If $\alpha^{\prime\prime}$ is disjoint from $\alpha^{\prime}$, then the proof is done. Otherwise they have bigons. At least one of these bigons bounds a topological disk and contains exactly one subarc of $\alpha^{\prime}$ and one subarc of $\alpha^{\prime\prime}$. By deforming isotopically the subarc of $\alpha^{\prime\prime}$ to the other side of $\alpha^{\prime}$, the curve $\alpha^{\prime\prime}$ can be deformed into an isotopic $\alpha_{1}$ with strictly less intersection points with $\alpha^{\prime}$ and also disjoint from $\alpha^{\prime\prime}$. Since the support of this deformation on $S$ is disjoint from $P$ and $\beta$, so $\alpha_{1}$ remains in minimal position with $o$ and is still disjoint from $\beta$. Hence we can conclude that $d^{\dagger}(\alpha_{1},o)=r$ and $d^{\dagger}(\alpha_{1},\alpha^{\prime\prime})=1$. Continuing the construction process, we gain $\alpha_{1},\ \alpha_{2},\ \dots,\ \alpha_{n}$ until $\alpha_{n}$ is disjoint from $\alpha^{\prime}$ in a way that $d^{\dagger}(\alpha_{i},o)=r$ and $d^{\dagger}(\alpha_{i},\alpha_{i+1})=1$ for all $11\leq i\leq n-$ 1. Hence we get a desired path $\alpha,\alpha^{\prime\prime},\alpha_{1},\dots,\alpha_{n},\alpha^{\prime}$. ∎
If we write $B_{r}(\gamma)\coloneqq\{\alpha\in\operatorname{\mathcal{NC}}^{\dagger}(S)\colon d^{\dagger}(\alpha,\gamma)\leq r\}$ the ball of radius $r$ centred at $\gamma$, then we will have the following result:
Lemma 4.5.
For every $\gamma\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ and $\alpha,\beta\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)\cap B_{1}(\gamma)$, then there exists a path
$\alpha=\nu_{0},\ \nu_{1},\ \dots,\ \nu_{\ell}=\beta$
in the graph $\operatorname{\mathcal{NC}}^{\dagger}(S)$ with
$\nu_{i}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)\cap B_{1}(\gamma)$
for $0\leq i\leq\ell$.
Proof.
Let $P\subset S$ be a non-empty finite subset so that $\{o,\gamma,\alpha\}$ are in minimal positions. If $\alpha$ and $\beta$ are isotopic in $S-P$, then by Lemma 4.4, we are done. Suppose now that they are not. Take a hyperbolic metric on $S-P$ so that $o$ and $\gamma$ become geodesic as per Proposition 3.4. Now we take the geodesic representatives $\alpha^{\prime}\in[\alpha]_{S-P}$ and $\beta^{\prime}\in[\beta]_{S-P}$. So $\{o,\gamma,\alpha^{\prime},\beta^{\prime}\}$ are pairwise in minimal position on $S-P$. In particular, we conclude that $\alpha^{\prime},\beta^{\prime}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)\cap B_{1}(\gamma)$ by Proposition 3.5. By sending the isotopy classes to their geodesic representatives on $S-P$, we can construct an isometric embedding $\iota\colon\operatorname{\mathcal{NC}}(S-P)\hookrightarrow\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$. We remark that [Wri23, Proposition 5.4] implies that there exists a path in $\operatorname{\mathcal{NC}}(S-P)$ connecting $[\alpha^{\prime}]_{S-P}$ to $[\beta^{\prime}]_{S-P}$ in which each vertices has distance $r+1$ to $[o]_{S-P}$ and distance $1$ to $[\gamma]_{S-P}$. Their images under $\iota$ then yield a path from $\alpha^{\prime}$ to $\beta^{\prime}$ included in $B_{1}(\gamma)$. Now we connect a path from $\alpha$ to $\alpha^{\prime}$ as well as from $\beta$ to $\beta^{\prime}$ as in Lemma 4.4. Because the vertices from these two paths are disjoint from $\gamma$, they also lie in $B_{1}(\gamma)$. By consequences, the concatenation of the three paths above yield the desired path. ∎
Now we have finished the proof of all assumptions in Proposition 2.11 for the subgraph $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)\subset\operatorname{\mathcal{NC}}^{\dagger}(S)\subset\mathcal{C}^{\dagger}(S)$. Note that they are all quasi-isometric. To prove Theorem 1.1, it suffices to apply Proposition 2.11 to $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$.
Proof of Theorem 1.1.
With $D=1$, for the subgraph $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$, Lemma 4.3 verifies the assumption (G1) and (G2) from Proposition 2.11 and Lemma 4.5 verifies the assumption (G3) from Proposition 2.11. ∎

### v4

@@ 5 Connectivity properties
5.A. Closed surfaces of higher genus.
Throughout this subsection, we will assume that $S$ is a closed surface of genus $g(S)\geq 3$. In this subsection, we will focus on verifying the conditions in Proposition 2.11 for the subgraph $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$.
For convenience, let $r>0$ be any positive integer, we denote by $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ the sphere in $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ centred at $o$ of radius $r$, i.e.
$\operatorname{\mathcal{S}_{o,\pitchfork}}(r)=\{\alpha\in\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)\colon d_{o}^{\dagger}(\alpha,o)=r\}.$
The following lemma shows that there is no “dead ends” in the sense of [birman2015curve] inside of the subgraph $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$, i.e. this subgraph satisfies the assumption (G1) in Proposition 2.11.
@@
Let $S$ be a closed hyperbolic surface. Let $\alpha\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ for some $r>0$. Then there exists a curve $\beta\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)$ such that $\beta\cap\alpha=\emptyset$.
@@
Let $o$ and $\alpha$ be as above and let $P\subset S$ be a finite collection of point such that $o$ and $\alpha$ are in minimal position of $S-P$. By Proposition 3.5, we can pick a hyperbolic metric on $S-P$ such that $o$ and $\alpha$ are geodesic.
Consider the subsurface $Y$ of $S-P$ given by cutting along $\alpha$. Note that the subsurface projection $\widehat{\pi_{Y}}\colon\mathcal{C}^{s}(S-P)\to\mathcal{C}(Y)$ is unbounded, as the subset of non-separating curves on $Y$, a fortiori surviving on $S-P$, is unbounded. We can take a non-separating geodesic curve $\beta\subset Y$ with respect to the hyperbolic metric on $S-P$ such that $d_{Y}\big{(}\widehat{\pi_{Y}}(o),\widehat{\pi_{Y}}(\beta)\big{)}>M$, where $M$ is the constant from Theorem 4.12. Since $\beta$ is non-separating on $S$, it is surviving on $S-P$. Now Theorem 4.12 implies that any geodesic in $\mathcal{C}^{s}(S-P)$ connecting $[o]_{S-P}$ to $[\beta]_{S-P}$ must contain an element that does not intersect $Y$ essentially. But as $\alpha$ is non-separating on $S$ (and a fortiori on $S-P$), the only curves that do not intersect $Y$ are those isotopic to $\alpha$. This implies that the geodesic in $\mathcal{C}^{s}(S-P)$ connecting $[o]_{S-P}$ to $[\beta]_{S-P}$ must pass through $[\alpha]_{S-P}$. As $\beta$ is disjoint from $\alpha$ by definition, we have
$d_{\mathcal{C}^{s}(S-P)}([o]_{S-P},[\beta]_{S-P})=d_{\mathcal{C}^{s}(S-P)}([o]_{S-P},[\alpha]_{S-P})+1.$ (5.1)
Because $o,\alpha,\beta$ are geodesic curves on $S-P$, they are pairwise in minimal position on $S-P$. Combining Proposition 3.7 and (5.1), we can conclude that $d^{\dagger}(o,\beta)=d^{\dagger}(o,\alpha)+1$ on $\mathcal{C}^{\dagger}(S)$. Since the inclusion map $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)\hookrightarrow\mathcal{C}^{\dagger}(S)$ is an isometric embedding (appealing to Lemma 3.12), we can conclude that $\beta\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)$. ∎
Now, we will treat in the rest of this subsection the condition (G3) in Proposition 2.11. First, we will need the following auxiliary lemmata:
@@
Let $r>0$ be any positive integer, and let $\alpha,\beta\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$. If $\alpha$ and $\beta$ are not transverse, then there exists a $\beta^{\prime}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ that is isotopic to $\beta$ yet disjoint from it, and which is also transverse to $\alpha$.
@@
In light of Proposition 3.7, we can find a finite subset $P\subset S$ so that $(o,\alpha)$ and $(o,\beta)$ are in minimal positions on $S-P$. By making a small perturbation as described in Lemma 3.9, we can obtain a curve $\beta^{\prime}$ that is transverse to $\alpha$ and $o$, disjoint from $\beta$, and isotopic to $\beta$ in $S-P$ while remains in minimal position with $o$ in $S-P$. With these assumptions, combining Proposition 3.7 and Lemma 3.12, we can conclude that
$d_{o}^{\dagger}(o,\beta^{\prime})=d^{\dagger}(o,\beta^{\prime})=d_{S-P}([o],[\beta^{\prime}])=d_{S-P}([o],[\beta])=d^{\dagger}(o,\beta)=d_{o}^{\dagger}(o,\beta)=r.$
This implies that $\beta^{\prime}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ and the proof is complete. ∎
Remark 5.3.
In the sequel, as for any $\alpha\in\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$, we have the distance equality between $d^{\dagger}(o,\alpha)$ and $d^{\dagger}_{o}(o,\alpha)$, we will no longer distinguish these two distances when working on the spheres.
Lemma 5.4.
Let $r>0$ and $\alpha,\widehat{\alpha}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ be isotopic on $S-P$, where $P$ is a finite subset of $S$. If $(o,\alpha)$ and $(o,\widehat{\alpha})$ are in minimal position on $S-P$, there exists a path on $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ connecting $\alpha$ to $\widehat{\alpha}$, in which every vertex is a curve isotopic to $\alpha$ on $S-P$. If in addition, if $\gamma\in\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ is disjoint from both $\alpha$ and $\widehat{\alpha}$, then the path can be chosen so that all vertices on the path are disjoint from $\gamma$.
Proof.
Up to a small perturbation as in Lemma 3.9, we can take a curve $\alpha_{1}$ that is isotopic $\alpha$ to but disjoint from $\alpha$ and transverse to $\widehat{\alpha}$ (and also to $o$). As the perturbation can be made sufficiently small, one can ensure that there is no bigon between the curve $\alpha_{1}$ and the curve $o$ on $S-P$, or equivalently they remain in minimal position. If $\alpha_{1}$ is also disjoint from $\widehat{\alpha}$, the proof is done. Otherwise, if $\alpha_{1}\cap\widehat{\alpha}\neq\emptyset$, there is at least one bigon between $\alpha_{1}$ and $\widehat{\alpha}$ as they are isotopic (so the minimal position is disjoint). Repeating the procedures of isotoping an subarc of $\alpha_{1}$ to the other side of $\alpha$, we can then construct a path $\alpha_{1},\alpha_{2},\dots,\alpha_{n}$ with $\alpha_{i}\cap\alpha_{i+1}=\emptyset$ for any $1\leq i<n$ and $\alpha_{n}\cap\widehat{\alpha}=\emptyset$. Moreover, as they remain in minimal position with $o$, by Proposition 3.7 and Lemma 3.12, we still have $d_{o}^{\dagger}(o,\alpha_{i})=r$ for all $i$, i.e. $\alpha_{1},\alpha_{2},\dots,\alpha_{n}$ form a desired path on $\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ connecting $\alpha$ to $\widehat{\alpha}$. For disjointness from $\gamma$, as $\gamma$ and $\alpha,\widehat{\alpha}$ contain a priori no bigon, at each step of the isotopy above, the newly constructed curve will not produce bigon with $\gamma$, i.e. they remain disjoint from $\gamma$. ∎
Lemma 5.4 indicates that under the assumption of condition (G2), i.e. two curves $\alpha,\beta\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)$ both adjacent to a curve $\gamma\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$, then we can connect $\alpha$ to $\alpha^{\prime}$ (respectively $\beta$ to $\beta^{\prime}$) by a path on $\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)$ where each vertex is again disjoint from $\gamma$, so that $o,\gamma,\alpha^{\prime},\beta^{\prime}$ are geodesic curves on a punctured surface $S$. In result, these curves all lie in an isometric copy of $\mathcal{C}^{s}(S-P)\hookrightarrow\mathcal{C}^{\dagger}(S)$, allowing us to deduce the condition (G2) with $D=2$ as [wright2023spheres].
To be precise, we will prove the following result:
Lemma 5.5.
Let $S$ be a closed surface of genus $g\geq 3$ and let $\gamma\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$ for some positive integer $r>0$ and $\alpha,\beta\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)$ that are adjacent to $\gamma$ in $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$. Then there exists a path
$\alpha=\nu_{0},\nu_{1},\dots,\nu_{\ell}=\beta$
in $\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)$ connecting $\alpha$ to $\beta$ with $\nu_{i}\cap\gamma=\emptyset$.
Proof.
Let $P\subset S$ be a finite collection of points such that the pairs $(o,\gamma)$, $(o,\alpha)$ and $(o,\beta)$ are in minimal position on $S-P$. Suppose that $S-P$ is carrying a hyperbolic metric such that $o$ and $\gamma$ are geodesic, of which the existence is guaranteed by Proposition 3.5. Let $\alpha^{\prime}$ and $\beta^{\prime}$ be the geodesic representative of $[\alpha]_{S-P}$ and $[\beta]_{S-P}$. By Lemma 5.4, we can already find paths on $\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)$ connecting $\alpha$ to $\alpha^{\prime}$ and $\beta$ to $\beta^{\prime}$ such that every vertices on the two paths are disjoint from $\gamma$. So to complete the proof, it suffices to prove the lemma for $\alpha^{\prime}$ and $\beta^{\prime}$.
Let $Y$ be the complementary component of $\gamma$ in $S-P$. Let $K>0$ be a positive integer, take
$\mathcal{O}(\gamma,K)\coloneq\left\{\nu\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)\colon d_{Y}\big{(}\widehat{\pi_{Y}}(o),\widehat{\pi_{Y}}(\nu)\big{)}>K\right\}.$
By appealing to the proof of [wright2023spheres, Lemma 7.2], we can show that for any $K>M$, where $M$ is the constant from Theorem 4.12, we can find a path connecting $\alpha^{\prime}$ to $\alpha^{\prime\prime}\in\mathcal{O}(\gamma,K)$ such that each vertices on the path is on $\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)$ and disjoint from $\gamma$. We remark that whenever the classical bounded geodesic image theorem is used in the proof of [wright2023spheres, Lemma 7.2], we will replace it by Theorem 4.12. Note that $Y$ has genus at least $2$ and $[\alpha]_{Y}$ is non-separating on $Y$. Inspecting the proof of [wright2023spheres, Lemma 7.4], the path above can be chosen such that all vertices are non-separating curves on $Y$ and a fortiori on $S$. The same can be done for $\beta^{\prime}$ connecting to a $\beta^{\prime\prime}\in\mathcal{O}(\gamma,K)$. Moreover, as $\alpha^{\prime\prime}$ and $\beta^{\prime\prime}$ are disjoint from $\gamma$, their $Y$-subsurface projection consists only their isotopy class on $Y$, i.e. $[\alpha^{\prime\prime}]_{Y},[\beta^{\prime\prime}]_{Y}\in\mathcal{C}(Y)$.
Now it remains to show that $\mathcal{O}(\gamma,K)$ is connected for some sufficiently large $K>M$. Use the connectedness of spheres in the curve graph of high complexity surfaces [wright2023spheres, Theorem 1.1], we can obtain a path in $\mathcal{C}(Y)$
$[\alpha^{\prime\prime}]_{Y}=q_{0},\dots,q_{n}=[\beta^{\prime\prime}]_{Y}$
such that $d_{Y}\big{(}q_{i},\widehat{\pi_{Y}}(o)\big{)}>K>M+C$, where $C$ is the constant in [wright2023spheres, Lemma 3.11] for the existence of a path connecting two non-separating curves on $Y$, within distance $C$ from the geodesic in $\mathcal{C}(Y)$ between theses two curves, consisting of only isotopy classes of non-separating curves on $Y$, i.e. [wright2023spheres, Lemma 3.11] implies that we can modified the above path into a new path
$[\alpha^{\prime\prime}]_{Y}=q_{0}^{\prime},\dots,q_{m}^{\prime}=[\beta^{\prime\prime}]_{Y}$
where $q_{i}^{\prime}$ are non-separating curves on $Y$ such that $d_{Y}\big{(}q_{i}^{\prime},\widehat{\pi_{Y}}(o)\big{)}>M$. By descending to their geodesic representatives on $S-P$, Theorem 4.12 applies to make it a path on $\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)$ connecting $\alpha^{\prime\prime}$ to $\beta^{\prime\prime}$ such that all vertices are disjoint from $\gamma$.
Now the concatenating path $\alpha\rightsquigarrow\alpha^{\prime}\rightsquigarrow\alpha^{\prime\prime}\rightsquigarrow\beta^{\prime\prime}\rightsquigarrow\beta^{\prime}\rightsquigarrow\beta$ will yield the desired path. ∎
Lemma 5.6.
Let $S$ be a closed surface of genus $g(S)\geq 3$. For every adjacent pair $\alpha,\beta\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$, there exists a path
$\alpha=\nu_{0},\ \nu_{1},\ \cdots,\ \nu_{\ell}=\beta$
such that $\nu_{i}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)$ for $0<i<\ell$ and $d^{\dagger}(\nu_{i},\alpha)\leq 2$.
Proof.
Let $P\subset S$ be a finite collection of points such that $o,\alpha,\beta$ are pairwise in minimal position.
  1. (Case 1)
    First suppose that $(\alpha\cup\beta)$ is either isotopic or non-separating on $S$. Then let $Y$ be the complementary component in $S-P$ of $\alpha\cup\beta$ that is not a punctured cylinder. Again, we can find a non-separating curve $\gamma$ on $Y$ such that $d_{Y}\big{(}\widehat{\pi_{Y}}(o),\widehat{\pi_{Y}}(\gamma)\big{)}>M$ for $M>0$ from Theorem 4.12. So any geodesic in $\mathcal{C}^{s}(S-P)$ connecting $[o]_{S-P}$ to $[\gamma]_{S-P}$ must contain an element that does not intersect $Y$ essentially, whereas the only isotopy classes of surviving curves that do not intersecting $Y$ essentially are $[\alpha]_{S-P}$ to $[\beta]_{S-P}$. Note that $\gamma$ is surviving on $S-P$ and disjoint from $\alpha,\beta$. By making $\gamma$ in minimal position with $o$ on $S-P$, similar to the case in Lemma 5.1, we can conclude that $\gamma\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)$.
  2. (Case 2)
    Now suppose that $(\alpha\cup\beta)$ is separating and let $Y_{1},Y_{2}\subset S-(\alpha\cup\beta)$ be two complementary components. As $g(S)\geq 3$, we observe that $Y_{1}$ has genus $g(Y_{1})\geq 1$. Similarly, take $\gamma\subset Y_{1}$ be a geodesic non-separating curve on $Y_{1}$ such that $d_{Y_{1}}\big{(}\pi_{Y_{1}}(o),[\gamma]_{Y_{1}}\big{)}>M$, where $M>0$ is the constant from Theorem 4.12. We claim that $d_{\mathcal{C}^{s}(S-P)}([o]_{S-P},[\gamma]_{S-P})\neq r-1$. Indeed, by Theorem 4.12, the geodesic in $\mathcal{C}^{s}(S-P)$ connecting $[o]_{S-P}$ to $[\gamma]_{S-P}$ must contain an element $[\xi]_{S-P}$ that dos not intersecting $Y_{1}$ essentially. If $d_{\mathcal{C}^{s}(S-P)}([o]_{S-P},[\gamma]_{S-P})=r-1$, then $d_{\mathcal{C}^{s}(S-P)}([o]_{S-P},[\gamma]_{S-P})\leq r-2$ and as a result, $[\xi]_{S-P}\neq[\alpha]_{S-P},[\beta]_{S-P}$ as $d_{\mathcal{C}^{s}(S-P)}([o]_{S-P},[\alpha]_{S-P})=d_{\mathcal{C}^{s}(S-P)}([o]_{S-P},[\beta]_{S-P})=r$. This forces that the geodesic representative $\xi$ to be contained in $Y_{2}$ and hence disjoint from $\alpha$ and $\beta$, making the concatenating path
    $[o]_{S-P}\rightsquigarrow[\xi]_{S-P}\text{ --- }[\alpha]_{S-P}$
    of length at most $r-1$. This contradicts our assumption that $\alpha\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$. Hence $d^{\dagger}(o,\gamma)=d_{\mathcal{C}^{s}(S-P)}([o]_{S-P},[\gamma]_{S-P})=r$ or $r+1$. If $d^{\dagger}(o,\gamma)=r+1$, then we have done the proof. If $d^{\dagger}(o,\gamma)=r$, then $(\alpha\cup\gamma)$ and $(\beta\cup\gamma)$ are non-separating and appealing to (Case 1), the two pairs yield $\nu_{1},\nu_{2}\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)$ both adjacent to $\gamma\in\operatorname{\mathcal{S}_{o,\pitchfork}}(r)$. Now applying Lemma (G2) to the triplet $(\nu_{1},\nu_{2},\gamma)$, we can get a path $\nu_{1}\rightsquigarrow\nu_{2}$ on $\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)$ such that all vertices are disjoint from $\gamma$. As a result, the concatenating path
    $\alpha\text{ --- }\nu_{1}\rightsquigarrow\nu_{2}\text{ --- }\beta$
    yields a path on $\operatorname{\mathcal{S}_{o,\pitchfork}}(r+1)$ connecting $\alpha$ to $\beta$ within $B_{\alpha}(2)$.
Combining the two cases together proves the desired result. ∎
Now we can deduce Theorem 1.1 for closed surfaces with higher genus.
@@
Indeed, $\partial\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ is connected and linear connected by Proposition 2.11 for closed surfaces with genus $g(S)\geq 3$, as Lemma 5.1, Lemma 5.5 and Lemma 5.6 apply to verifies respectively the conditions (G1), (G2) and (G3). Note that by Proposition 3.11, $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ is quasi-isometric to $\operatorname{\mathcal{NC}}^{\dagger}(S)$ and as a result $\partial\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)\simeq\partial\operatorname{\mathcal{NC}}^{\dagger}(S)$. Finally, we remark that for a closed surface with genus $\geq 3$, we know that $\operatorname{\mathcal{NC}}^{\dagger}(S)$ is quasi-isometric to $\mathcal{C}^{\dagger}(S)$ with $\partial\operatorname{\mathcal{NC}}^{\dagger}(S)\simeq\mathcal{C}^{\dagger}(S)$ (see [bowden2022quasi]). This concludes the desired result. ∎